40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7903     Entry time: Wed Jan 16 08:29:45 2013
Author: Steve 
Type: Update 
Category: SUS 
Subject: PRM damping restored 

PRM oplev servo turned off.  OLPIT servo gain 0.15 and OLYAW  -0.3 set to ZERO.  PRM damping restored

ELOG V3.1.3-